# Conditional normality and finite-state dimensions revisited

### v1

@@ 1 Introduction
Let us start by recalling what is (unconditional) normality and how it is related to finite-state gambling. (For the details, proofs and references to original papers see, e.g., [2].)
@@ 1.1 Block frequencies
[…]
As we have noted, we have $\dim(\alpha)=1$ (and therefore $\operatorname{\mathrm{Dim}}(\alpha)=1$) for normal sequence $\alpha$. One can prove that the reverse is also true: if $\dim(\alpha)=1$, then $\alpha$ is normal ($\operatorname{\mathrm{Dim}}(\alpha)=1$ is not enough).
Instead of splitting $\alpha$ into $k$-bit blocks, we may consider a sliding window of size $k$, and consider the distribution of $k$-bit strings among the first $N$ non-aligned blocks (thus using only $N+k-1$ bits of the sequence instead of $Nk$). Then we proceed as before; one can prove that the finite-state dimension and strong finite-state dimension remain the same (as for splitting into aligned blocks).
@@ 1.2 Gambling and gales
The finite-state dimensions of a bit sequence $\alpha$ have natural interpretation in terms of gambling against $\alpha$. Consider a guessing game where a gambler starts with initial capital $1$ and, before getting the next bit of $\alpha$, makes a bet by splitting her capital into bets on 00 and $1$; one part is lost and the other is doubled. In general, the gambler strategy can be fully described by a function $m$ defined on binary strings: $m(X)$ is the capital of the gambler after having played against binary string $X$. The gambling rules mean that this function is a martingale, i.e., $m(\varepsilon)=1$ for the empty string $\varepsilon$, and $m(X)=[m(X0)+m(X1)]/2$ for every bit string $X$ and its two possible extensions $X0$ and $X1$. We restrict our attention to finite-state martingales that correspond to gambling strategies with finite memory. This means that at every moment gambler is in one of finitely many states; the state determines the ratio of the bets on 00 and $1$ (we assume the ratio is a rational number); the next state is determined by the previous one and the observed bit.
Intuitively, if a sequence $\alpha$ is “non-random”, the regularities in $\alpha$ can be used to gamble against it; more non-randomness means faster growth of the martingale along the sequence. In the ultimate case when the sequence is all zeros and we bet on zero all the time, the capital after $n$ bets is $2^{n}$. On the other hand, a cautious gambler that always splits the capital equally between 00 and $1$ keeps her original capital untouched (the martingale is always $1$). We measure the growth in logarithmic scale and consider the ratio $(\log m(a_{1}\ldots a_{n}))/n$. In the two extreme cases mentioned this ratio is $1$ and 00 respectively. In the general case, for a given sequence $\alpha=a_{1}a_{2}\ldots$ and finite-state gambler $G$ (with corresponding martingale $m_{G}$), we consider
@@
[…]
Usually this procedure is explained in a slightly different language: instead of measuring the capital growth rate, we impose a “tax” by multiplying the capital by $2^{s-1}$ after each game (a zero tax if $s=1$, and $50\%$ tax for $s=0$; in the latter case we need to put all the capital on the correct bit all the time just to keep the capital unchanged). In other words, we consider finite-state $s$-gales that satisfy the equation
@@
One can consider also combinations of finitely many martingales ($l$-account gales). This means that the initial capital is split between $l$ finite-state gamblers and each of them plays independently using her own capital. Note that while the total memory of $l$ gamblers is finite, they cannot be replaced by one finite-state gambler (since the combined ratio depends on the current capitals of $l$ individual gamblers and does not necessarily belong to any finite set). Still this generalization does not affect the finite-state dimension and strong finite-state dimension. (For finite-state dimension it is easier to see than for strong finite-state dimension, but it is true in both cases.)
@@ 1.3 Introducing oracles (conditions)
Informally speaking, conditional randomness of a bit sequence $\alpha=a_{1}a_{2}\ldots$ with respect to some other bit sequence $\beta=b_{1}b_{2}\ldots$ (considered as a condition) means that we cannot find any regularities in $\alpha$, or cannot win the gambling game against $\alpha$ even if we are given access to $\beta$ as an oracle (for free). For most notions of algorithmic randomness the conditional versions are well understood (and follow the standard relativization scheme used in computability theory: all computable objects are replaced by $\beta$-computable objects, where $\beta$ is used as an oracle). However, for finite-state randomness this general scheme does not work: it is not clear a priori how the finite memory gambler should access the oracle. If we think about this more closely, we see that there are some important choices — at least two of them — that should be made when we define conditional normality.
@@ 1.3.1 Look-ahead while gambling
But do we permit looking ahead or not? Assume that $\alpha=a_{1}a_{2}\ldots$ is some “perfectly random” sequence and $\beta=b_{1}b_{2}\ldots$ is $\alpha$ shifted to the left: $b_{n}=a_{n+1}$; does this make $\alpha$ non-normal with respect to $\beta$? The answer suggested by Nandakumar, Pulari and S in [1] is yes, we do allow looking ahead for $O(1)$ future oracle bits, and $\alpha$ in this example is not $\beta$-normal. More precisely, for each finite-state gambler some look-ahead constant $c$ is fixed; the gambler sees next $c$ bits $b_{i+1}\ldots b_{i+c}$ of the oracle $\beta$ (as well as $b_{i}$ itself) when gambling on some bit $a_{i}$ of $\alpha$.
@@ 1.3.2 Negligible frequencies
What do we require from this sequence of pairs? The second components (being part of an oracle that we do not control) can be arbitrary. What we could require is that the conditional distribution on first components, for every fixed value of the second component, is uniform. Note that this conditional distribution may be undefined if some string $B$ never appears as a condition (as the second component of a pair), or can be rather irrelevant if $B$ appears only finitely many times (say, once or twice) as a second component.
@@ 1.4 What follows
It turns out that the two approaches to relative normality described above give the same notion — this is one of the main results of [1]. (Note that the block approach is consistent with the permission to look ahead: condition block is provided in its entirety, not bit by bit.) Moreover, it is shown in [1] that these two approaches lead to equivalent definitions of conditional finite-state dimension and conditional strong finite-state dimension. The proof in [1] is quite technical, and in this note we show how these results can be obtained in a natural way by extending other definitions of finite-state dimensions (with finite-state complexity, superadditive functions etc.) to the conditional case and by proving equivalence of all these extended definitions.
@@ 2.2 Finite-state gamblers
Consider a game with infinite number of rounds. Initially the gambler has capital $1$. In $i$th round, before seeing $a_{i}$, the gambler sees $b_{i},\ldots,b_{i+c}$ (where $c$ is a constant that is chosen by the gambler before starting the game), and splits her current capital into two parts labeled 00 and $1$ (the ratio is a rational number chosen by the gambler). Then $a_{i}$ is shown to the gambler, the part labeled $a_{i}$ is doubled and the other part is lost.
@@
For every gambling strategy $G$ (finite-state or not) we consider the corresponding martingale function whose value $m_{G}(a_{1}\ldots a_{N}\mskip 0.8mu|\mskip 0.8mub_{1}\ldots b_{N+c})$ is the gambler capital after playing against $a_{1}\ldots a_{N}$ with oracle bits $b_{1}\ldots b_{N+c}$ according to the rules. Then we define dimensions as
@@ 2.3 Automatic complexity
Here $|P|$ stands for the length of the binary string $P$ and $D$ is a ternary relation on binary strings $A,B,P$ called a description mode; we read $(A,B,P)\in D$ as “$P$ is a description of $A$ given $B$”. In Kolmogorov complexity theory we require that $D$ is a function, namely, the first argument $A$ is a function of two other arguments $B$ and $P$, and this function should be computable. For every $D$ that has these properties we consider the corresponding function $\operatorname{\mathrm{C}}_{D}$. There exist optimal description modes $D$ that make $\operatorname{\mathrm{C}}_{D}$ minimal up to $O(1)$ additive term; we fix one of them and call $\operatorname{\mathrm{C}}_{D}(A\mskip 0.8mu|\mskip 0.8muB)$ the (plain) Kolmogorov complexity of $A$ given $B$.
@@ 2.4 Superadditive complexity measures
Remarks.
1. The first condition somehow reflects the finite-state requirement; it says, roughly speaking, that no information transfer happens between two stages where $A_{1}$ and $A_{2}$ are described. (It is technically important that we do not allow any increase and have no $O(1)$ term in the inequality). As we will see, for every automatic description mode the function $\operatorname{\mathrm{C}}_{D}$ is superadditive. Note also that for the standard Kolmogorov complexity this requirement does not hold: $K(A_{1}A_{2}\mskip 0.8mu|\mskip 0.8muB_{1}B_{2})$ could be much smaller than $K(A_{1}\mskip 0.8mu|\mskip 0.8muB_{1})+K(A_{2}\mskip 0.8mu|\mskip 0.8muB_{2})$ if $A_{1}$ and $A_{2}$ share a lot of information.
2. The second requirement says that $K$ should not be “too small” (e.g., the zero function is not a superadditive complexity measure even though it is superadditive). It is fulfilled for conditional Kolmogorov complexity and for automatic complexity since (for a given condition) each string $P$ may be description of only one (for Kolmogorov complexity) or $O(1)$ (for automatic complexity) strings.
@@ 2.5 Finite-state a priori complexity
Now we modify this definition to adapt it to the finite-state case (taking into account the synchronous access to bits of condition/oracle). Consider a graph with finitely many vertices (states). For every state, and for every condition bit (00 or $1$), there are two outgoing edges that correspond to output bits 00 and $1$, labeled with rational non-negative numbers with sum $1$. (In total for every vertex we have two pairs with sum $1$: one for condition bit 00 and one for condition bit $1$, each of these four numbers is written on the edge going to the next state after the transition). These numbers are interpreted as probabilities to emit 00 and $1$ (and change the state according to their edges) being in a given state and reading a given condition (input) bit.
@@
Then, as in the Kolmogorov complexity theory, we define the finite-state a priori complexity:
@@
Remark.
In this definition we do not allow the random process to arrive to different states but output the same bit (for a given input bit). This is natural if we think about connection with finite-state gamblers (the probabilities to emit 00 and $1$ correspond to the parts of the gambler’s capital bet on 00 and $1$), but a more general definition can also be used (see the remark at the end of Section 3.5).
@@ 3.1 From block frequencies to automatic complexity
[…]
@@
  • Finally, the Shannon–Fano theorem has some overhead: the average length of code (per block) is bounded by $H+1$, not $H$. This overhead ($+1$) then will be divided by block size $k$, but even $1/k$ overhead is bad for us: we need exact equality.
How do we deal with these problems? For the first one, we use not one prefix-free code but a family of prefix-free codes provided by the following lemma.
@@
so $H(A^{1}\mskip 0.8mu|\mskip 0.8muB^{1})+H(A^{2}\mskip 0.8mu|\mskip 0.8muB^{2})<2\tau$, as we promised. So we may switch to larger blocks without increase in entropy (per bit), and the $O(1)$ overhead becomes twice less important. This construction then can be repeated (or we may switch directly from $k$ to any multiple $k^{\prime}$ of $k$).
[…]
@@
[…]

### v2

@@ 1 Introduction
Let us start by recalling what (unconditional) Borel normality is and how it is related to finite-state gambling. (For the details, proofs, and references to original papers see, e.g., [2].)
@@ 1.1 Block frequencies
More general, we consider the entropy $H_{k,N}(\alpha)$ of the distribution on the first $N$ blocks of length $k$. Then we consider $\liminf$ and $\limsup$ of this entropy as $N$ goes to infinity. For normal sequences the distribution converges to the uniform distribution, so both $\liminf$ and $\limsup$ are $k$ (for every $k$). This gives a characterization of normal sequences. In a general case we divide $\limsup_{N}$ and $\liminf_{N}$ of $H_{k,N}(\alpha)$ by $k$ and take the limit as $k\to\infty$ (one can prove that this limit always exists and coincides with the infimum over $k$). These two limits (for $\liminf$ and $\limsup$) are called respectively finite-state dimension of $\alpha$, denoted by $\dim_{\mathrm{FS}}(\alpha)$, and strong finite-state dimension of $\alpha$, denoted by $\operatorname{\mathrm{Dim}}_{\mathrm{FS}}(\alpha)$. The letters “FS” stand for “finite-state”; we omit them in the sequel since we do not consider other dimensions.
As we have noted, we have $\dim(\alpha)=1$ (and therefore $\operatorname{\mathrm{Dim}}(\alpha)=1$) for a normal sequence $\alpha$. One can prove that the reverse is also true: if $\dim(\alpha)=1$, then $\alpha$ is normal ($\operatorname{\mathrm{Dim}}(\alpha)=1$ is not enough).
Instead of splitting $\alpha$ into $k$-bit blocks, we may consider a sliding window of size $k$, and consider the distribution of $k$-bit strings among the first $N$ non-aligned blocks (thus using only N+k1N+k-1 bits of the sequence instead of $Nk$). Then we proceed as before; one can prove that the finite-state dimension and strong finite-state dimension remain the same (as for splitting into aligned blocks).
@@ 1.2 Gambling and gales
The finite-state dimensions of a bit sequence $\alpha$ have natural interpretation in terms of gambling against $\alpha$. Consider a guessing game where a gambler starts with initial capital $1$ and, before getting the next bit of $\alpha$, makes a bet by splitting her capital into bets on 0 and $1$; one part is lost and the other is doubled. In general, the gambler strategy can be fully described by a function $m$ defined on binary strings: $m(X)$ is the capital of the gambler after having played against binary string $X$. The gambling rules mean that this function is a martingale, i.e., $m(\varepsilon)=1$ for the empty string $\varepsilon$, and $m(X)=[m(X0)+m(X1)]/2$ for every binary string $X$ and its two possible extensions $X0$ and $X1$. We restrict our attention to finite-state martingales that correspond to gambling strategies with finite memory. This means that at every moment gambler is in one of finitely many states; the state determines the ratio of the bets on 0 and $1$ (we assume that this ratio is a rational number); the next state is determined by the previous one and the observed bit.
Intuitively, if a sequence $\alpha$ is “non-random”, the regularities in $\alpha$ can be used to gamble against it; more non-randomness means faster growth of the martingale along the sequence. In the ultimate case when the sequence is all zeros and we bet on zero all the time, the capital after $n$ bets is $2^{n}$. On the other hand, a cautious gambler that always splits the capital evenly between 0 and $1$ keeps her original capital untouched (the martingale is always $1$). We measure the growth in logarithmic scale and consider the ratio $(\log m(a_{1}\ldots a_{n}))/n$. In the two extreme cases mentioned this ratio is $1$ and 0 respectively. In the general case, for a given sequence $\alpha=a_{1}a_{2}\ldots$ and a finite-state gambler $G$ (with corresponding martingale $m_{G}$), we consider
@@
Then we take supremum over finite-state gamblers $G$ (we are interested in the best gamblers, not in the worst ones) and get two quantities that are directly related to finite-state dimension: one can prove that the first one (with $\limsup$) is $1-\dim(\alpha)$, and the second one (with $\liminf$) is $1-\operatorname{\mathrm{Dim}}(\alpha)$.
Usually this procedure is explained in a slightly different language: instead of measuring the capital growth rate, we impose a “tax” by multiplying the capital by $2^{s-1}$ after each game (a zero tax if $s=1$, and $50\%$ tax for $s=0$; in the latter case we need to bet all the capital on the correct bit all the time just to keep the capital unchanged). In other words, we consider finite-state $s$-gales that satisfy the equation
@@
One can consider also combinations of finitely many martingales ($l$-account gales). This means that the initial capital is split between $l$ finite-state gamblers and each of them plays independently using her own capital. Note that while the total memory of $l$ gamblers is finite, they cannot be always replaced by one finite-state gambler, since the combined ratio depends on the current capitals of $l$ individual gamblers and does not necessarily belong to any finite set. Still this generalization does not affect the finite-state dimension and strong finite-state dimension. (For finite-state dimension it is easier to see than for strong finite-state dimension, but it is true in both cases.)
@@ 1.3 Introducing oracles (conditions)
Informally speaking, conditional randomness of a bit sequence $\alpha=a_{1}a_{2}\ldots$ with respect to some other bit sequence $\beta=b_{1}b_{2}\ldots$ (considered as a condition) means that we cannot find any regularities in $\alpha$, or cannot win the gambling game against $\alpha$ even if we are given access to $\beta$ as an oracle (for free). For most notions of algorithmic randomness the conditional versions are well understood (and follow the standard relativization scheme used in computability theory: all computable objects are replaced by $\beta$-computable objects, where $\beta$ is used as an oracle). However, for finite-state randomness (normality) this general scheme does not work: it is not clear a priori how the finite memory gambler should access the oracle. If we think about this more closely, we see that there are some important choices — at least two of them — that should be made when we define conditional normality.
@@ 1.3.1 Look-ahead while gambling
But do we permit looking ahead or not? Assume that $\alpha=a_{1}a_{2}\ldots$ is some “perfectly random” sequence and $\beta=b_{1}b_{2}\ldots$ is $\alpha$ shifted to the right: $a_{n}=b_{n+1}$; does this make $\alpha$ non-normal with respect to $\beta$? The answer suggested by Nandakumar, Pulari and S in [1] is yes, we do allow looking ahead for $O(1)$ future oracle bits, and $\alpha$ in this example is not $\beta$-normal. More precisely, for each finite-state gambler some look-ahead constant $c$ is fixed; the gambler sees next $c$ bits $b_{i+1}\ldots b_{i+c}$ of the oracle $\beta$ (as well as $b_{i}$ itself) when gambling on some bit $a_{i}$ of $\alpha$.
@@ 1.3.2 Negligible frequencies
What do we require from this sequence of pairs? The second components (being part of an oracle that we do not control) can be arbitrary. What we could require is that the conditional distribution on first components, for every fixed value of the second component, is uniform. Note that this conditional distribution may be undefined if some string $B$ never appears as a condition (as the second component of a pair), or can be rather irrelevant if $B$ appears only finitely many times as a second component.
@@ 1.4 What follows
It turns out that the two approaches to relative normality described above give the same notion — this is one of the main results of [1]. (Note that the block approach is consistent with the permission to look ahead: condition block is provided in its entirety, not bit by bit.) Moreover, it is shown there that these two approaches lead to equivalent definitions of conditional finite-state dimension and conditional strong finite-state dimension. The proof in [1] is quite technical, and in this note we show how these results can be obtained in a natural way by extending other definitions of finite-state dimensions (with finite-state complexity, superadditive functions etc.) to the conditional case and by proving equivalence of all these extended definitions (thus anwering also some questions left open in [1]).
@@ 2.2 Finite-state gamblers
Consider a game with infinite number of rounds. Initially the gambler has capital $1$. In $i$th round, before seeing $a_{i}$, the gambler sees $b_{i},\ldots,b_{i+c}$ (where $c$ is a constant that is chosen by the gambler before starting the game), and splits her current capital into two parts labeled 0 and $1$ (the ratio is a rational number chosen by the gambler). Then $a_{i}$ is shown to the gambler, the part labeled $a_{i}$ is doubled and the other part is lost.
@@
For every gambling strategy $G$ (finite-state or not) we consider the corresponding martingale function whose value $m_{G}(a_{1}\ldots a_{N}\mskip 0.8mu|\mskip 0.8mub_{1}\ldots b_{N+c})$ is the gambler’s capital after playing against $a_{1}\ldots a_{N}$ with oracle bits $b_{1}\ldots b_{N+c}$ according to the rules. Then we define dimensions as
@@ 2.3 Automatic complexity
Here $|P|$ stands for the length of the binary string $P$ and $D$ is a ternary relation on binary strings $A,B,P$ called a description mode; we read $(A,B,P)\in D$ as “$P$ is a description of $A$ given BB”. In Kolmogorov complexity theory we require that $D$ is a function, namely, the first argument $A$ is a function of two other arguments $B$ and $P$, and this function should be computable; in other words, $D$ should be (computably) enumerable. For every $D$ that has these properties we consider the corresponding function $\operatorname{\mathrm{C}}_{D}$. There exist optimal description modes $D$ that make $\operatorname{\mathrm{C}}_{D}$ minimal up to $O(1)$ additive term; we fix one of them and call $\operatorname{\mathrm{C}}_{D}(A\mskip 0.8mu|\mskip 0.8muB)$ the (plain) Kolmogorov complexity of $A$ given $B$.
@@
Remark 2.
The connection between different notions of compressibility and entropy goes back to Shannon, and the relation between finite-state compression and block entropy was analyzed long ago in [5] where the notion of finite-state (strong) dimension appeared (under the name of compressiblity and denoted by $\rho$). See [2, Section 6] for the historic account of these and subsequent developments. What seems to be new in our approach (used in [2] for non-conditional complexity and adapted here for conditional complexity) is that we consider only decompression (as it is done in algorithmic information theory) and choose the technical details in a special way to guarantee superadditivity that plays an important technical role in the proofs.
@@ 2.4 Superadditive complexity measures
Remark 3.
The first condition somehow reflects the finite-state requirement; it says, roughly speaking, that no information transfer happens between two stages where $A_{1}$ and $A_{2}$ are described. (It is technically important that we do not allow any increase and have no $O(1)$ term in the inequality). As we will see, for every automatic description mode the function $\operatorname{\mathrm{C}}_{D}$ is superadditive. Note also that for the standard Kolmogorov complexity this requirement does not hold: $K(A_{1}A_{2}\mskip 0.8mu|\mskip 0.8muB_{1}B_{2})$ could be much smaller than $K(A_{1}\mskip 0.8mu|\mskip 0.8muB_{1})+K(A_{2}\mskip 0.8mu|\mskip 0.8muB_{2})$ if $A_{1}$ and $A_{2}$ share a lot of information. (The formula for the complexity of pairs guarantees subadditivity with logarithmic precision.)
Remark 4.
The second requirement says that $K$ should not be “too small” (e.g., the zero function is not a superadditive complexity measure even though it is superadditive). It is fulfilled for conditional Kolmogorov complexity and for automatic complexity since (for a given condition) each string $P$ may be description of only one (for Kolmogorov complexity) or $O(1)$ (for automatic complexity) strings.
@@ 2.5 Finite-state a priori complexity
Now we modify this definition to adapt it to the finite-state case taking into account the synchronous access to bits of condition/oracle. Consider a graph with finitely many vertices (states). For every state, and for every condition bit (0 or $1$), there are two outgoing edges that correspond to output bits 0 and $1$, labeled with rational non-negative numbers with sum $1$. In total for every vertex we have two pairs with sum $1$: one for condition bit 0 and one for condition bit $1$, each of these four numbers is written on the edge going to the next state after the transition. These numbers are interpreted as probabilities to emit 0 and $1$ (and change the state according to their edges) being in a given state and reading a given condition (input) bit.
@@
Then, as in the algorithmic information theory, we define the finite-state a priori complexity:
@@
Remark 5.
In this definition we do not allow the random process to arrive to different states but output the same bit (for a given input bit). This is natural if we think about connection with finite-state gamblers (the probabilities to emit 0 and $1$ correspond to the parts of the gambler’s capital bet on 0 and $1$), but a more general definition can also be used (see the remark at the end of Section 3.5).
@@ 3.1 From block frequencies to automatic complexity
We start by proving that the block entropy dimensions cannot be smaller than the corresponding automatic complexity dimensions. In other words, we assume that for some $k$ the limit ($\liminf$ or $\limsup$) of conditional $k$-bit block entropies is smaller than some threshold $\tau$, and we construct an automatic description mode for which the corresponding limit is also smaller than $\tau$.
@@
  • Finally, the Shannon–Fano theorem has some overhead: the average length of code (per block) is bounded by $H+1$, not $H$. This overhead (+1+1) then will be divided by block size $k$, but even $1/k$ overhead is bad for us: we need exact equality.
How do we deal with these problems? [^1] In [2], unfortunately, this is done incorrectly for the case of strong dimensions (sorry!): Lemma 19.1 is not enough since it provides prefix code for $N$-bit strings that depends on $N$, and this cannot be used to construct one finite automaton for all $N$. We need to be more careful and construct a finite family of codes for strings of fixed length, as it is done in Lemma Lemma and Lemma Lemma. For the first one, we use not one prefix-free code but a family of prefix-free codes provided by the following lemma.
@@
so $H(A^{1}\mskip 0.8mu|\mskip 0.8muB^{1})+H(A^{2}\mskip 0.8mu|\mskip 0.8muB^{2})<2\tau$, as we promised. So we may switch to larger blocks without increase in entropy (per bit), and the $O(1)$ overhead becomes twice less important. This construction then can be repeated (or we may switch directly from $k$ to any multiple of $k$).
This almost finishes the proof; there is only one subtle point that we missed. The doubling argument assumes that the value of $N$ where the block entropy was small, is even (then we may compare the $k$-block entropy with $2k$-block entropy). However, we are not guaranteed that the values of $N$ that are important for $\liminf$, are even. To overcome this difficulty, we note that entropy function (for a given block size) is uniformly continuous on a simplex where it is defined, and adding/deleting one block for large $N$ causes only small change in the frequencies, so we may add or delete one block and still have entropy smaller than $\tau$. (In fact, for $\limsup$ we do not have this problem, since in this case entropy is smaller than $\tau$ for all large $N$, including large even values of $N$.)
@@
and the same is true for $\limsup$ instead of $\liminf$ (in both places).